OpenStudy (anonymous):

The function f has the property that f(x)=f(x+1) for all numbers x. If f(4)=17, what is the value of f(8)?

OpenStudy (whpalmer4):

If we know that \(f(x) = f(x+1)\) for all values of \(x\), then if we know \(f(4)=17\), what does \(f(5)=\)

OpenStudy (anonymous):

um is it 6?

imqwerty (imqwerty):

To get a feel for this information, let's first pick a random value for x. Let's say x = 2 The given info tells us that f(x) = f(x + 1) for all values of x So, f(2) = f(2 + 1) In other words, f(2) = f(3) Let's try a different value of x. How about x = 3 The given info tells us that f(x) = f(x + 1) for all values of x So, f(3) = f(3 + 1) In other words, f(3) = f(4) And so on. We can see that f(2) = f(3) = f(4) = f(5) and so on So, if f(4) = 17, we can be certain that f(8) = 17 as well.

OpenStudy (anonymous):

wait why is f(8)=17? I'm so confused ....

OpenStudy (whpalmer4):

\[f(x) = f(x+1)\]is just a way of saying that whatever the value of the function \(f(x)\) is at some value of \(x\), it has the same value at \(x+1\)

OpenStudy (anonymous):

yes i understand that but y is f(8)=17?

OpenStudy (whpalmer4):

we know that \(f(4) = 17\) right?

OpenStudy (anonymous):

yes

imqwerty (imqwerty):

Another approach is to create some kind of function that satisfies the condition that f(x) = f(x + 1) for all values of x, AND also f(4) = 17 Here's one: f(x) = (x)(0) + 17 Here we can see that: f(1) = 17 f(2) = 17 f(2) = 17 f(3) = 17 f(4) = 17 . . . f(8) = 17

OpenStudy (whpalmer4):

okay, and we know that whatever value \(f(x)\) has, \(f(x+1)\) has the same value

OpenStudy (anonymous):

yes

OpenStudy (whpalmer4):

so that means that \(f(4) = f(4+1) = f(5)\) and we already know that \(f(4) = 17\) so that means \(f(5) = 17\) as well

OpenStudy (anonymous):

wait what? why is f(5)=17 now?

OpenStudy (whpalmer4):

for any value of \(x\), pick a number, any number, \(f(x) = f(x+1)\) or \(f(x+1) = f(x)\)

OpenStudy (whpalmer4):

the problem tells us that..."for all numbers x"...

OpenStudy (anonymous):

okay yes

OpenStudy (whpalmer4):

okay, so if \{f(4) = 17\), then we are taking \(x = 4\) and we can write \(f(4) = f(4+1) = 17\) because \(f(x) = f(x+1)\)

OpenStudy (anonymous):

yes

OpenStudy (whpalmer4):

okay, if f(5) = 17, what does f(6)? remember, 5+1 = 6

OpenStudy (anonymous):

wait.... so this is a trick question? everything equals to 17?

imqwerty (imqwerty):

yes it is tricky

imqwerty (imqwerty):

yomamafb u r forgetting somthing

OpenStudy (anonymous):

?

OpenStudy (anonymous):

oh i just saw your reply ..... so EVERYTHING equals to 17? wth so this was a trick question!

imqwerty (imqwerty):

u were going to tell a number

OpenStudy (whpalmer4):

yes, exactly, in this case, the value of \(f(x) = 17\) for every value of \(x\)

OpenStudy (anonymous):

fml

hero (hero):

f(2) = f(3) f(3) = f(4) f(4) = f(5) ...

OpenStudy (anonymous):

it was a trick question everything equals 17